A Porsche 944 Turbo has a rated engine power of 217 hp. 30% of the power is lost in the drive train, and 70% reaches the wheels. The total mass of the car and driver is 1420 kg, and two-thirds of the weight is over the drive wheels.
(a) What is the maximum acceleration of the Porsche on a concrete surface where
Answer #1
P = FV
P= 0.70 Pmax
Fmax V = 0.70 Pmax
V = 0.70 Pmax / Fmax

f = mus (2/3)mg
f = Fmax
F=ma (and Fmax)
ma = mus (2/3) mg
amax= mus (2/3) g
amax =  0.92 * (2/3) * 9.81
= 6.0 m/s2

------------------------------------------

V = 0.70 Pmax / Fmax
V = 0.70 Pmax/ mamax
V = 0.70 x 217 x 746/ (1420 x 6.0)

V=13.3 m/s
-------------------------------------------
t = V/a= 13.3 /6.0
= 2.22 s